OpenStudy (jemson):

From a group of boys and girls, 15 girls leave. There are then left 2 boys for each girl. After this, 45 boys leave. There are then 5 girls for each boy. The number of girls in the begining was???

OpenStudy (lxelle):

the key is inwriting these things as equations. (b=boys, g=girls) "15 girls leave. Two boys remain for each girl." means b/(g-15) = 2 then, we need 2 equations for 2 unknowns, so we use the other bit. "45 boys leave and 5 girls remain for each boy." (g-15)/(b-45) = 5 Now, I just rearrange the first equation to say (g-15)=b/2 then I plug that in the second equation. (b/2)/(b-45)=5 solve for b b=10b-450 450=9b b=50 and if b=50, we can plug that into the first equation. (g-15)=(50)/2 g-15=25 g=??

OpenStudy (triciaal):

@LXelle did you check your answer?

OpenStudy (lxelle):

Yeah, I guess I did.

OpenStudy (lxelle):

Anything to add in?

OpenStudy (lxelle):

After you got the g, plus both g and b together then , you'll get your final answer. :))

OpenStudy (triciaal):

I did not work through the whole problem I had G - 15 = 2B ; B - 45= G/5 then your response popped up and I went through

OpenStudy (triciaal):

b/(g-15) = 2 this would mean 2 girls for every boy

OpenStudy (anonymous):

Call B the number of boys and G the number of girls at the beginning. 15 girls leave and you have 2 boys for each girl, so: B = 2 ( G-15) Then, 45 boys leave and you have 5 girls for each boy, so: 5 (B- 45) = G -15 Hence you have a system of linear equations: B - 2 G = - 30 5 B - G = 210 and the solution is that at the beginning there were B= 50 boys and G = 40 girls in the group.